OpenStudy (anonymous):

y=sin(x+pi/4) Find the zero(s) of f on the interval [0, 2π). (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.)

OpenStudy (anonymous):

\[\sin(0)=0\] so you need \[x+\frac{\pi}{4}=0\] solve for x get \[x=-\frac{\pi}{4}\] but of course that is not in your interval, so add \(\pi\) and get \[x=\frac{3\pi}{4}\]

OpenStudy (anonymous):

x+pi/4=pi*k where k in an integer

OpenStudy (anonymous):

then add \(\pi\) again and get \[x=\frac{7\pi}{4}\]

OpenStudy (anonymous):

oh thank so much=)

OpenStudy (anonymous):

why do i add pi, not 2pi?